MIDC News
Year of 2006
December 27, 2006
- 11:38 MST: November 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 11:28 MST: November 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 17:29 MST: Due additional security restrictions, which will take place in upcoming MIDC server migration, the current SRRL SkyCam images will be updated every 5 minutes (instead of on-demand). Note: Internal NREL users will still have access to the instantanous snapshots. -Afshin Andreas
- 09:33 MST: Datacollection to PFCI has resumed. Data was lost from Oct 31 @ 13:38 MST to Nov 10 @ 12:10 MST. -Afshin Andreas
- 15:25 MST: September & October 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 15:19 MST: September & October 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 08:58 MST: There is a communications problem with the PFCI datalogger. Last collection occured on October 31, 2006 @ 5:36 MST.
-Afshin Andreas - 08:57 MST: As of 24-OCT-2006, The Phoenix Federal Correctional Institution (PFCI) data is now available on the MIDC. The station has a Rotating Shadowband Pyranometer (RSP), a wind set, and temp/RH. Data starts on 1-FEB-2006, however the RSP was not fully functional until 29-SEP-2006. -Afshin Andreas
- 11:26 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service @ 8:21 MST, calibrated against NIST F407, and returned to service @ 11:14 MST today. -Afshin Andreas
- 10:58 MST: The SRRL BMS Upwelling CM3 (on CNR1) was calibrated against the BMS Upwelling PSP using 7/22/06 data. New cal factor was entered into the logger @ 9:57 MST today. -Afshin Andreas
- 12:09 MST: All SRRL BMS Global & Diffuse UV instruments were calibrated on 9-12-2006, this includes: UV-S-B-T, CUVB1, MS210W, UVB-1 (global & diffuse), 501A, UV-S-A-T, CUVA1, and TUVR (global). New cal factors were entered into the logger @ 12:00 MST today. -Afshin Andreas
- 13:01 MST: August 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 12:51 MST: August 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 10:53 MST: The SRRL BMS Direct UVA and UVB sensors were calibrated on 9-5-2006. New cal factors were entered into the logger today @ 10:49 MST. -Afshin Andreas
- 07:54 MST: There was an unknown communcations problem with the SRRL ATI over the weekend. Data was lost from Aug 19 @ 7:06 through Aug 20 @ 1:13 MST. -Afshin Andreas
- 14:09 MST: The UNLV update frequency has been changed from 3 hours to 60 minutes. -Afshin Andreas
- 14:50 MST: July 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 14:29 MST: July 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 06:56 MST: The SRRL BMS UV-S-A-T & UV-S-B-T were swapped back to their normal position at 06:46 MST today. -Afshin Andreas
- 15:30 MST: The SRRL BMS UV-S-A-T and UV-S-B-T intruments were swapped at 15:28 MST today for an over-night instrument test (cal factors were NOT swapped in logger). Instruments will be restored tommorow morning. -Afshin Andreas
- 07:39 MST: July 20th 2006 0745MST - M2 PSP ventilator repaired and replaced. -Pete Gotseff
- 12:48 MST: The SRRL BMS spare Direct RG780 NIP was calibated in BORCAL 2006-02. Instrument was swapped out 12:35-12:43 MST and logger was updated at 12:44 MST today. -Afshin Andreas
- 11:55 MST: The SRRL BMS Direct CUVA2 was mistakenly disconnected at the junction box. Unit is operational as of 11:06 MST today. -Afshin Andreas
- 13:27 MST: The old SRRL BMS Global 90-North LI-200 S/N PY1724 was put back into service and logger updated at 13:24 MST. A replacement LI-200 will be put into the next BORCAL 2006-04 which will replace PY1724. -Afshin Andreas
- 11:36 MST: There is a problem with the SRRL BMS Global 90-North LI-200 S/N PY1726, that was installed yesterday. It seems to be reading too high. The calibration history shows a continous decline in response from 15 µV/W/m2 in 1998 to 9.7 today. Instrument will likely be discarded and replaced in the near future. -Afshin Andreas
- 08:25 MST: The NWTC M2 Global PSP was swapped out with an instrument that was calibrated in BORCAL 2006-01, between 7:51 and 8:04 MST today. Ventilator was removed temporarily due to bird contamination. Datalogger was updated with new cal factor @ 8:19 MST. -Afshin Andreas
- 14:23 MST: The SRRL BMS Global CM3 (mounted on the CNR1 at radiometer tower) was calibrated using BORCAL 2006-01 reference global data. New cal factor were entered into the logger at 13:34 MST today. -Afshin Andreas
- 14:22 MST: The following SRRL BMS radiometer tower spare instruments were calibrated in BORCAL 2006-02: Globals 90-West (PSP & LI-200), Global 90-North (PSP & LI-200), Global 40-South (PSP & LI-200). Instruments were swapped out between 13:00 to 13:42 MST today. New cal factors were entered into the logger at 13:34 MST today. -Afshin Andreas
- 08:41 MST: The heater in the SRRL BMS CNR1 was enabled for the first time @ 8:24 MST today. Hopefully this will dry the water drops on the CG3 instrument (flat dome), which cause very large measurement errors in the downwelling infrared. -Afshin Andreas
- 13:25 MST: June 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 13:18 MST: June 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 13:48 MST: The SRRL BSRN spare Diffuse PSP and Direct NIP was calibated in BORCAL 2006-02. Instruments were swapped out 13:26-13:28 MST and logger was updated at 13:30 MST today. -Afshin Andreas
- 07:20 MST: The SRRL BMS Direct TUVR was shorted at 7:13 MST today, in an effort to "reset" it. Signal has been dropping out since Feb 24 2006. -Afshin Andreas
- 09:49 MST: There is still a problem with the SRRL BMS Direct CUVA2, unit is reading about 45 times low (and polarity is now reversed). -Afshin Andreas
- 13:13 MST: The SRRL BMS Direct CUVA2 was returned to service today @ 13:04 MST. The unit was repaired at the factory and will need calibration sometime soon. -Afshin Andreas
- 12:14 MST: The following SRRL BMS instruments were calibrated in BORCAL 2006-02: Global CM22, CM6b, and LI-200. New factors were entered into the logger @ 11:55 MST today. -Afshin Andreas
- 11:09 MST: The SRRL SkyCam web camera was replaced with a recently acquired used "mint condition" camera (same make and model) at 11:00 MST today. The old jerry-rigged camera had mismatched electronics, which caused wierd sunny images. -Afshin Andreas
- 08:54 MST: There is a bug in the SRRL AOCS software that causes it to not produce a spectra within 10 minutes on either side of solar noon (when it should be). -Afshin Andreas
- 11:07 MST: Communication with the SRRL Global Horiz LI-1800 was re-established today @ 10:45 MST. The LICOR TransTerm was used to "reset" the instrument. -Afshin Andreas
- 07:03 MST: Communication with the SRRL Global Horiz LI-1800 was lost this morning. Problem unknown at this time. -Afshin Andreas
- 12:34 MST: The SRRL Global Horiz LI-1800 data since 10:25 MST has been moved to it's own data stream. A new "global horizontal" selection is now available from the Spectral Data page. -Afshin Andreas
- 10:54 MST: The SRRL Global 40-South LI-1800 spectroradiometer was removed from service @ 9:15 MST and re-installed in a Global Horizontal position at 10:25 MST. The data stream is currently in label as "40-South," it will be moved to it's own "Horizontal" data stream later today or this week. -Afshin Andreas
- 13:37 MST: The following SRRL BMS spare instruments were calibrated in BORCAL 2006-01: Diffuse 8-48, Diffuse PSP, Diffuse PSP (shadowband), and Global RG780. Instruments were swapped out between 13:04 to 13:10 MST today. New cal factors were entered into the logger at 13:10 MST today. -Afshin Andreas
- 14:47 MST: The SRRL BMS instruments mentioned in the June 6 @ 13:15 MST News Bullet, were updated with the corrected factors at 14:01 MST. -Afshin Andreas
- 14:41 MST: As of 14:01 MST today, the SRRL BMS data stream has been upgraded to v10.2. All historical data since Jan. 1, 2006 has been converted to the new format. The following changes occured:
- Global PSP (ventilated/zenith corrected) is now using Effective Net IR Corrected Coefficients to calculate irradiance.
- Global PSP (non-ventilated) using Effective Net IR Corrected Coefficients to calculate irradiance has been added.
- Diffuse PSP (ventilated/thermal corrected) is now using Effective Net IR Corrected Cal Factor @ 45o to calculate irradinace. (Note: The CFc@45 in use is currently estimated, instrument will be replaced with calibrated instrument tomorrow)
- (4) Reserved Channels have been added for the CR23X data logger (Research 6,7,8,9).
- 09:45 MST: The SRRL BSRN spare Global PSP was calibated in BORCAL 2006-01. Instrument was swapped out and logger updated at 9:24 MST today. -Afshin Andreas
- 09:44 MST: The following SRRL BMS radiometer tower spare instruments were calibrated in BORCAL 2006-01: Globals 90-East (PSP & LI-200), Global 90-South (PSP & LI-200), Upwelling (PSP & LI-200), and Global Normal CM-21. Instruments were swapped out between 8:20 to 8:51 MST today. New cal factors were entered into the logger at 8:43 MST today. -Afshin Andreas
- 14:29 MST: May 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 14:25 MST: May 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 13:35 MST: The SRRL BMS instrument cal factors that were updated on May 31 (see 08:17 MST news item) are using the wrong factor, because the BORCAL 2006-01 data was reprocessed using a new Cavity WRR factor. The change in cal factor is very small and will be corrected in the datalogger sometime this month (during a datastream upgrade for Pyranometer Effective Net IR Correction). -Afshin Andreas
- 15:07 MST: The SRRL BSRN NIP tracker has been removed from service, due to problems that started yesterday @ ~9:30 MST. -Afshin Andreas
- 08:17 MST: The following SRRL BMS instruments were calibrated in BORCAL 2006-01: Globals: PSP, PSP (vent), PYR-P, TSP-1, TSP-700, SP-Lite; Directs: CH1, NIP #1, NIP #2, LI-201; Diffuse: CM-22. New cal factors and PSP (vent) function were entered into the logger & data acquisition system at 7:48 MST. -Afshin Andreas
- 14:58 MST: Due to installation of new equipment lift to the SRRL deck, power was shutdown on 5/25/2006 from 10:50 to 11:10 MST. All diffuse pyranometers and pyrgeometers were unshaded. All pyheliometers and photometers were not pointing at the sun. All UV instruments (except Eppley TUVRs), all Yankee pyranometers, and the AOCS were down. All ventilators for pyranometers were OFF. The appropriate Data Quality Statements (DQS) for affected instruments have been updated. -Afshin Andreas
- 10:20 MST: March & April 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 09:09 MST: March & April 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 14:25 MST: Increased the logger range for the SRRL BMS Photometric LI-210 sensor at 14:22 MST, which would over-range occasionaly. -Afshin Andreas
- 12:29 MST: The NWTC M2 instrument page was updated for all instruments that were changed on April 20, this includes all wind speed, all wind direction (except 2m), all temperatures, and the dew point sensor. -Afshin Andreas
- 15:52 MST: New calibration constants for the NWTC M2 W/S and Temp sensors were entered into the logger program at 15:48 MST today. -Afshin Andreas
- 08:21 MST: The NWTC M2 tower was dropped at 7:58 MST today for instrument changeout. -Afshin Andreas
- 15:05 MST: Datacollection for the SRRL TSI-880, Schott RSP, and ATI RSP resumed today @ 14:50 MST. All data for RSPs is up to date. -Afshin Andreas
- 07:13 MST: Datacollection for the SRRL ATI & Schott RSP is currently down, since yesterday at 13:00 MST, due to phone line re-routing. -Afshin Andreas
- 14:28 MST: The SRRL TSI-880 was disconnected at 13:00 MST today, for ethernet re-routing. Cloud Cover is temporarily unavailable. -Afshin Andreas
- 12:42 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service @ 9:41 MST, calibrated against NIST F407, and returned to service @ 12:19 MST today. -Afshin Andreas
- 14:37 MST: The SRRL BMS upwelling quantum sensor is now functional, it was replaced with S/N Q20555 at 14:33 MST. -Afshin Andreas
- 16:24 MST: The SRRL BMS Upwelling Quantum sensor was removed from service @ 12:38 MST today. -Afshin Andreas
- 07:46 MST: There is a problem with the SRRL BMS Upwelling Quantum Sensor since March 23 at approx 9:30 MST. The sensor is reading low and/or negative. -Afshin Andreas
- 15:32 MST: February 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 15:16 MST: February 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:54 MST: The SRRL BMS Global PSP was returned to service @ 09:07 MST today. The original Global Ventilated PSP was swapped back into service (with original CF & coeff.) @ 09:28 MST today. -Afshin Andreas
- 08:14 MST: The SRRL BMS Diffuse PSP was returned to service @ 07:57 MST today. The Global PSP was removed from service @ 07:58 MST for characterization in the Black Body System. -Afshin Andreas
- 11:51 MST: The following SRRL BMS IR data (irradiance and instrument nets) was corrected to use the new K-factors that were generated on 2/24/2006:
- DW CG3 010284: Feb 1 @ 0:00 through Feb 24 @ 15:01 MST
- UW CG3 010284: Feb 1 @ 0:00 through Feb 24 @ 15:58 MST
- 09:26 MST: The SRRL BMS Global Ventilated PSP was replaced with S/N 25823F3 @ 9:07 MST today (CF & Coeffs. also updated), so that the BMS instrument's RSnet could be characterized in the NREL Black Body System. The SRRL BMS Diffuse Ventilated PSP was also removed for the same reason @ 9:16 MST, it's output is set to missing (-99999). -Afshin Andreas
- 16:05 MST: Corrected a coding problem in the SRRL BMS Rad Tower Logger @ 15:59 MST for the Upwelling CG3 caused by NREL/SGP BB Coef. Eqn. update. -Afshin Andreas
- 16:02 MST: Updated SRRL BMS Radiometer Tower Logger to use new NREL/SGP BB Coefficients Equation for the DW & UW CG3s at 15:02 MST. This also required updated the K-factors for these instruments. -Afshin Andreas
- 09:15 MST: The SRRL BMS Downwelling CG4 data (irradiance and instrument net) was corrected to use the new K-factors that were generated on 2/23/2006 from 2/1/2006 @ 0:00 MST to 2/23/2006 @ 12:51 MST. -Afshin Andreas
- 16:46 MST: Records for the SRRL ATI RSP & Schott RSP have been added to the Data Quality Statements (DQS) database. See their instrument history pages for links to the database. -Afshin Andreas
- 12:54 MST: The K1 and K2 factors for the SRRL BMS CG4 were corrected in the logger today @ 12:52 MST. -Afshin Andreas
- 11:05 MST: The SRRL BMS Downwelling PIR data (irradiance and instrument net) was corrected to use the new K-factors that were generated on 2/22/2006 from 2/1/2006 @ 0:00 MST to 2/22/2006 @ 10:10 MST. -Afshin Andreas
- 11:03 MST: The K0 & K2 factors for the SRRL BMS Downwelling PIR were corrected in the logger program on Feb 22 @ 10:11 MST. -Afshin Andreas
- 11:45 MST: The SRRL Schott RSP data from Feb 17 @ 4:00 MST through Feb 20 @ 12:00 MST is bogus due to a discharged battery. -Afshin Andreas
- 14:16 MST: New K-factors for the SRRL BMS Downwelling PIR were entered into the logger program today @ 14:08 MST. -Afshin Andreas
- 14:49 MST: Updated SRRL RSP Instrument History Page to include information about all output values. -Afshin Andreas
- 09:10 MST: January 2006 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 08:20 MST: The original ATI head unit (including LICOR PY21094) was returned to service on Jan 30 @ 11:19 MST. -Afshin Andreas
- 14:54 MST: Updated MIDC Monthly Report Generator to include "percent untested" in the "data recovery" information box at the bottom of reports. Reports for January 2006 and forward will reflect this update. -Afshin Andreas
- 14:44 MST: January 2006 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:26 MST: The following SRRL BMS IR data (irradiance and instrument nets) was corrected to use the new K-factors that were generated on 1/30/2006:
- DW CG4 010548: Jan 11 @ 14:17 through Jan 30 @ 9:04 MST
- UW PIR 31192F3: Jan 12 @ 12:00 through Jan 30 @ 9:04 MST
- UW CG3 010284: Jan 20 @ 11:17 through Jan 30 @ 10:32 MST
- 10:35 MST: The SRRL BMS Upwelling CG3 calibration constants were corrected in the logger program today @ 10:33 MST. -Afshin Andreas
- 09:22 MST: The SRRL BMS Upwelling PIR and Downwelling CG4 calibration constants were corrected in the logger program today @ 9:05 MST. -Afshin Andreas
- 11:41 MST: The SRRL BMS Upwelling CG3 calibration constants were updated in the logger program today @ 11:17 MST. -Afshin Andreas
- 13:02 MST: The 1-MIN Peak Gust @ 42' has been added to the SRRL BMS instantaneous download/display (including the graphical display). This is the peak wind speed that was recorded during the last minute (and throughout the current minute), based on the 1-second samples read by the data logger. The graphical display now uses the gust value (instead of the "current" wind speed) in the summary table. -Afshin Andreas
- 11:30 MST: Some electronic components in the SRRL SkyCam were swapped out today around 11:00 MST in an attempt to diagnose the problem and jerry-rig a solution. Images may look bogus. -Afshin Andreas
- 11:24 MST: The SRRL ATI head unit (including the LICOR PY21094) was swapped out with a unit to be calibrated (S/N PY17942) on 1/16/2006 @ 10:49 MST. The data is currently out of calibration. -Afshin Andreas
- 07:53 MST: The problem with the SRRL SkyCam is due to a problem with the webcam, which was most likely damaged during the cold spell in Dec 2005. Current and historical images since then will likely be bogus or unavailable. -Afshin Andreas
- 12:41 MST: For calibration purposes, the SRRL BMS data stream was upgraded to v10.1 at 12:00 MST today, the following Instrument Nets were added: PIR (UpWelling), CG3 (DownWelling), & CG3 (UpWelling). All data since Jan. 1, 2006 has been converted to the new format. -Afshin Andreas
- 15:10 MST: For calibration purposes the SRRL BMS upwelling and downwelling PIRs were swapped between ~14:00 to ~14:30 MST today. Datalogger program was updated at 14:17 MST for the swap out and updated K1 factors for the Upwelling PIR, Downwelling CG3 & CG4. -Afshin Andreas
- 15:53 MST: The Yankee TSI-880 Total sky images and "cloud analyzed" images (with cloud cover values), which are stored in the YESDAQ database can now be accessed through MIDC. Images are available since July 14, 2005 at 10 minute intervals. -Afshin Andreas
- 15:48 MST: Changed MIDC plotting behavior of missing values, such that missing values are ignored (not plotted), instead of set to zero. Note: If all data on a particular y-axis is missing (and auto scale is on) then the following error will result: "Error: Plot output is empty, there may be no valid data." -Afshin Andreas
- 09:44 MST: The SRRL BMS Direct CUVA2 was removed from service Jan. 5 @ 11:25 MST for troubleshooting. -Afshin Andreas
- 14:34 MST: December 2005 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 14:26 MST: December 2005 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:04 MST: The SRRL BMS datastream v9 to v10 change-over-point has been moved from Dec. 21, 2005 to Jan. 1, 2006. This required converting data from v10 to v9 format during the end of December. Any new v10 instruments were moved to BMS research channels during this time. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023]